Let f(x)=3x-3,g(x)=x+1,h(x)=3x+12. Then (fogohogohogohogoh)(0)=


1
Expert's answer
2021-05-07T10:23:40-0400

"h(x)=3x+12\newline
g(h(x))=g(3x+12)=3x+13\newline
h(g(h(x)))=h(3x+13)=9x+51\newline
g(h(g(h(x))))=g(9x+51)=9x+52\newline
h(g(h(g(h(x)))))=h(9x+52)=27x+168\newline
g(h(g(h(g(h(x))))))=g(27x+168)=27x+169\newline
h(g(h(g(h(g(h(x)))))))=h(27x+169)=81x+519\newline
g(h(g(h(g(h(g(h(x))))))))=g(81x+519)=81x+520\newline
f(g(h(g(h(g(h(g(h(x)))))))))=f(81x+520)=243x+1557\newline
\text{put x=0},\newline
f(g(h(g(h(g(h(g(h(0)))))))))=1557"